Problem: The students from algebra class go for coffee together every day at 9:00 AM. On Monday, Jordan paid $5.45 for three cartons of milk, four cups of coffee, and seven doughnuts. On Tuesday, Ryan paid $5.30 for four milks, two coffees, and eight doughnuts. On Wednesday, Zack paid $5.15 for two milks, five coffees, and six doughnuts.
(1) Set up a system of equations to model this scenario.
(2) Determine the cost of one carton milk, one coffee, and one doughnut by solving the system.
(3) If on Thursday, Taylor had to pay for five milks, two coffees and nine doughnuts, then how much did he spend?

Answer by mangopeeler07(462)

On Monday, Jordan paid $5.45 for three cartons of milk, four cups of coffee, and seven doughnuts. On Tuesday, Ryan paid $5.30 for four milks, two coffees, and eight doughnuts. On Wednesday, Zack paid $5.15 for two milks, five coffees, and six doughnuts.

(1) Set up a system of equations to model this scenario.
m=price milk d=price doughnuts c=price coffee

3m+4c+7d=5.45
4m+2c+8d=5.30
2m+5c+6d=5.15

This is the system you want to solve. Or you can solve:

3m+4c+7d=545
4m+2c+8d=530
2m+5c+6d=515

The only difference is that you would get your answers in terms of cents, not dollars.